OpenStudy (anonymous):

Find the equation of each line in slope-intercept form: Perpendicular to the line y=-2x+5 through the point(8,-3)

OpenStudy (anonymous):

IMPORTANT LINE RELATED EQUATIONS TO KNOW AND MEMORIZE slope formula m= slope/ gradiant -- same thing \[m=\frac{y_2-y_1}{x_2-x_1}\] standard formula \[Ax+By=C\] point-slope formula \[y-y_1=m(x-x_1)\] slope-intercept formula b= y-intercept -- in the form of (0,y) \[y=mx+b\]

OpenStudy (anonymous):

perpendicular slopes are negative reciprocals to each other

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

A perpendicular slope is going to be one where the original slope times the new slope = -1, so since your initial slope is -2, your new slope is going to be 1/2 because (-2)(1/2) = -1. Good so far?

OpenStudy (anonymous):

Next, we do point-slope: y - (-3) = (1/2)(x - 8) Next, slope-intercept, which is what we want: y + 3 = (1/2)x - 4 y = (1/2)x - 7 And that's done! Is that understandable?

OpenStudy (anonymous):

good

OpenStudy (anonymous):

ok f(x) \[F(x)= x ^{2} + 3, find f (-4)\]

OpenStudy (anonymous):

Is that F(x) = x^2 + 3 ? It came out jumbled on my screen.

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

find f if f = 4

OpenStudy (anonymous):

-4 i mean

OpenStudy (anonymous):

You just replace the "x" with "-4" -> (-4)^2 + 3 -> 16 + 3 -> 19